OpenStudy (anonymous):

The number of real roots of (3-x)^4 + (5-x)^4 = 16 is?

OpenStudy (anonymous):

Expand the first side and let what you get =16

OpenStudy (anonymous):

the answer should be 2...

OpenStudy (anonymous):

\[ (3-x)^4=x^4-12 x^3+54 x^2-108 x+81\\ (5-x)^4 =x^4-20 x^3+150 x^2-500 x+625\\ \] Now add both and set equal to 16

OpenStudy (anonymous):

Actually x=3 is a solution.

OpenStudy (anonymous):

and x=5 is also a solution.

OpenStudy (anonymous):

There are two real solutions 3 and 5 and two complex solutions\[ 4 \pm i \sqrt 7 \]

OpenStudy (anonymous):

\[ (3-x)^4+(5-x)^4-16=\\2 x^4-32 x^3+204 x^2-608 x+690=\\2 (x-5) (x-3) \left(x^2-8 x+23\right) \] Since 5 and 3 were easily found, you can divide by (x-5)(x-3) and get the factorization above.

OpenStudy (anonymous):

The answer to the problem is 2.